1.what is the length of the diagomnal of a rectangle whose length is 16 and whose width is 12?
2.If the legs of a right triangle measure 3 and 4 units,what is the measure of the hypotenuse?
3.The hypotenuse of a right triangle measures 13 and one leg measures 12. what is the measure of the other leg?

Answer by checkley79(3341)

#1 12^2+16^2=C^2
144+256=C^2
400=C^2
C=SQRT400
C=20 ANS.
#######################
#2
A^2+B^2=C^2
3^2+4^2=C^2
9+16=C^2
25=C^2
C=SQRT25
C=5 ANS.
#####################
YOU SHOULD BE ABLE TO SOLVE THE NEXT ONE.